# Cost-Free Personalization via Information-Geometric Projection in Bayesian Federated Learning

### v1

@@ Abstract
Bayesian Federated Learning (BFL) combines uncertainty modeling with decentralized training, enabling the development of personalized and reliable models under data heterogeneity and privacy constraints. Existing approaches typically rely on Markov Chain Monte Carlo (MCMC) sampling or variational inference, often incorporating personalization mechanisms to better adapt to local data distributions. In this work, we propose an information-geometric projection framework for personalization in parametric BFL. By projecting the global model onto a neighborhood of the user’s local model, our method enables a tunable trade-off between global generalization and local specialization. Under mild assumptions, we show that this projection step is equivalent to computing a barycenter on the statistical manifold, allowing us to derive closed-form solutions and achieve cost-free personalization. We apply the proposed approach to a variational learning setup using the Improved Variational Online Newton (IVON) optimizer and extend its application to general aggregation schemes in BFL. Empirical evaluations under heterogeneous data distributions confirm that our method effectively balances global and local performance with minimal computational overhead.
@@ 1 Introduction
Federated learning (FL) is a collaborative machine learning paradigm designed to preserve data privacy. By connecting a central server to multiple participants, commonly referred to as clients or end-users, FL enables distributed training while keeping data local to the clients. In general, FL algorithms operate by training local models on the private datasets of clients and aggregating these models into a global model maintained on the server. After performing local updates, clients share their updated model parameters with the server, which aggregates them to refine the global model. The updated global model is then sent back to the clients for further local training [37].
@@
In real-world scenarios, statistical heterogeneity among clients is a frequent challenge. For example, variations in MRI machines across hospitals can result in feature skewness, where input data distributions differ significantly between clients. Similarly, label distribution shifts can arise when some clients have disproportionately more samples from certain classes compared to others. This often occurs in specialized hospitals dedicated to specific diseases, as opposed to general hospitals. Such heterogeneity violates the theoretical assumption of independent and identically distributed (i.i.d.) data, leading to significant performance degradation and increased uncertainty in the global model when evaluated on local test data.
@@
Personalized Federated Learning (PFL) is a paradigm in which clients can adapt the global model to their local data, i.e., model personalization, with the objective of limiting the impact of statistical heterogeneity. In [45], the authors propose a distinctive taxonomy for PFL approaches, categorizing them into global model personalization techniques, e.g., client selection [49], meta-learning [18] and transfer-learning [9], and learning personalized models techniques, e.g., parameter decoupling [4], knowledge distillation [26] and clustering [42]. Furthermore, multiple studies [51, 6, 32, 52] applied Bayesian methods to the PFL framework, achieving better model calibration and uncertainty quantification. The transition from deterministic FL to Bayesian FL (BFL) naturally leads to an investigation of the manifolds to which clients’ and global posteriors distributions belong, allowing for the exploration of their properties and the operations that can be defined on them, as suggested in [25].
@@
In this work, we study the PFL paradigm through the lens of Bayesian learning, leveraging Information Geometry to develop a personalization method for BFL. Unlike existing approaches, our method does not require fine-tuning, additional training, or access to local or shared global data, making it a fully private and truly cost-free personalization technique. By specifying a divergence metric, we learn the personalized model posteriors by projecting the global model posterior distribution onto a sphere centered at the local model posterior distribution, where the radius encodes the desired degree of personalization the end-user would like to achieve. This method provides unlimited personalization flexibility, empowering end-users to seamlessly control the trade-off between local and global performance at no additional cost.
We show that for any divergence that is convex in its first argument, the projection problem is equivalent to computing the weighted barycenter of the local posterior and the global posterior distributions. This connection between two key concepts in Information Geometry, namely, geometric projection and the geometric barycenter, offers an interpretable link between the radius of the sphere and the weights for the barycentric aggregation. As a result, the degree of personalization becomes more intuitive and accessible to end-users.
The paper is structured as follows. Section 2 provides an extensive analysis of the related work. In Section 3, we present the parametric BFL framework based on variational learning for local training and posterior aggregation to determine the posterior global model. We propose a design that extends the prior use [40] of the Improved Variational Online Newton optimizer (IVON) [43] for various aggregation methods. In Section 4, we introduce the novel idea we present to personalize global BFL models using information-geometric projection and show its equivalence to barycentric aggregation for divergence functions convex in their first argument. Worth noting that the personalization method is not limited to parametric BFL, as the idea of personalization remains general and valid for the different settings of BFL, non-parametric included. Subsequently, we refine our results for the parametric case where, for specific divergences, we can leverage closed-form solutions for the barycenters, leading to substantial gains in algorithmic efficiency. We dedicate Section 5 to the experimental analysis, where we compare various aggregation techniques under the variational learning setting. Furthermore, we evaluate our method against state-of-the-art approaches and demonstrate that it achieves a favorable global/local trade-off compared to existing baselines.
2 Background and Related Work
Bayesian Federated Learning
BFL seeks to integrate the advantages of Bayesian Deep Learning, e.g., better model calibration and uncertainty quantification, within the FL framework. Following the classification proposed in [7], BFL approaches can be broadly categorized into two main types; Client-side BFL [51, 52, 36, 6, 23, 5] and server-side BFL [10, 8, 1, 22], with certain overlaps between these groups, reflecting the diversity in model formulations and inference techniques. In [25], the authors extend the discussion on the parametric client-side BFL framework by categorizing methods according to the techniques used to aggregate the local posteriors and propose a unified framework for aggregation rooted in Information Geometry.
However, for the scope of this work, it is useful to introduce an additional categorization of BFL based on the adopted modeling assumptions, distinguishing between parametric and nonparametric Bayesian methods.
  • Parametric BFL: Parametric Bayesian learning constitutes a principled framework within Bayesian statistics, in which the data generation process is assumed to adhere to a parametric model. Specifically, this approach assumes that the data distribution can be fully characterized by a finite set of parameters. The parametric family most commonly considered in BFL is the Gaussian family, where both client-specific posteriors and the global model posterior are approximated using Gaussian distributions [51, 39, 22, 29, 40, 44]. This Gaussian assumption facilitates tractable inference and efficient parameter aggregation.
  • Nonparametric BFL: Unlike parametric BL, non-parametric Bayesian learning doesn’t restrict the model to a specific family of distributions. Instead, it provides the flexibility to learn directly from the data, allowing the structure and complexity to emerge naturally rather than being predefined. This approach leverages distributions over infinite-dimensional function spaces, such as Gaussian Processes or Dirichlet Processes employed in the context of BFL in [50]. Additionally, Particle-Based VI, utilized in [27] for BFL, aligns with the non-parametric perspective by using a set of particles to approximate posteriors without assuming a predefined parametric family.
Personalized Bayesian Federated Learning
Personalized BFL extends the standard BFL framework by incorporating client-specific adaptations to the shared global model. FedPop [32] captures this personalization by modeling each client’s data generation process as a combination of fixed shared population parameters, describing the common data model, and client-specific random effects, describing the heterogeneity in the clients’ data, thus providing personalization and uncertainty estimation capabilities. To infer these parameters efficiently, FedPop employs MCMC methods to approximate local posterior distributions and perform stochastic optimization in a federated setting. Although MCMC ensures flexible and asymptotically exact Bayesian inference, it can be computationally expensive in large-scale settings. To address this, pFedBayes [51] employs variational inference to approximate client-specific posteriors, optimizing a tractable Evidence Lower Bound (ELBO) to achieve computational efficiency in BFL. Although pFedBayes presents a promising approach to personalization in BFL, certain aspects of its methodology introduce notable limitations. Firstly, the degree of personalization is uniformly applied across all clients and determined empirically, without an adaptive mechanism or an intuitive explanation to tailor it to individual client characteristics. Secondly, it defines the local model as personalized based on the use of VI, where the global model serves as a prior within the KL divergence term of the ELBO. This formulation may be controversial within the Bayesian community, as the classical interpretation of a prior belief should be independent of the observed data. More recently, pFedVEM [52] has taken advantage of variational inference to estimate the specific uncertainty of the client and the deviation of the model, modeling the client parameters as Gaussian distributed around a global latent variable. It employs a confidence-based aggregation strategy that ensures that clients with lower uncertainty and fewer deviations contribute more to the global model.
Unlike the aforementioned methods, our approach presents personalization as an additional step independent from the training process, as it only requires access to the global and local posteriors with no need for extra data or fine-tuning. In addition, our approach allows access to three different variants of the model, i.e., local, personalized, and global models, as suggested in [16], with the possibility of continually adjusting the personalized variant without additional training cost.
Information Geometry and Optimal Transport in FL
Information Geometry [3] provides a general framework for the study of the geometric properties of statistical manifolds when considering divergence functions, allowing us to intuitively treat distributions as elements in a manifold. This framework allows for the natural translation of fundamental geometric concepts, such as projections onto constraint sets [12, 13] and the identification of barycenters, serving as statistical centers of mass for a given set of probability distributions [38, 17].
On the other hand, Optimal Transport [46] focuses on identifying the most efficient transport plan between two probability distributions in terms of a given cost function. Under specific assumptions on the cost function, optimal transport allows one to define distance functions on the statistical manifold, as shown by the family of Wasserstein-p distances [46]. The Sinkhorn algorithm [14] serves as a natural bridge between Information Geometry and Optimal Transport, allowing for efficient computation of the transport plan by incorporating entropy regularization into the Optimal Transport problem.
Multi-Marginal Optimal Transport (MMOT) [21, 41] extends the classical two-marginal Optimal Transport problem to multiple distributions by seeking an optimal joint coupling that minimizes a given $n$-ary cost function of all marginal distributions. Furthermore, for specific cases of the $n$-ary cost, it can be shown that the MMOT problem is equivalent to the computation of the Wasserstein Barycenter (WB) of the set of marginals. This equivalence highlights the MMOT as a more general framework, within which the WB emerges as a special case, characterizing the optimal interpolation of multiple distributions in the Wasserstein space. Applications of concepts from the Information Geometry and Optimal Transport literature have recently been explored in the FL setting. In [19], the authors introduce FedOT, a PFL algorithm that integrates Optimal Transport with model training. FedOT employs deterministic FL, leveraging MMOT to align data distributions. Specifically, it learns transport maps to transform data points from heterogeneous distributions into a shared domain while simultaneously training a predictive model on the mapped data. In a similar approach, FedDRO [34] aggregates client data distributions via a Wasserstein barycenter and trains the global model against worst-case perturbations within a Wasserstein ball around this barycenter. In the context of BFL, [24] considers structured latent variable models by keeping local latent variables private, formulating a decentralized variational inference problem for which a communication-efficient aggregation based on Wasserstein barycenters is proposed. In [25], the authors propose a unifying framework for aggregation in BFL based on barycentric aggregation.
3 Parametric Bayesian Federated Learning
3.1 Learning Phase
The primary objective of BFL is to estimate the posterior distribution of the global model parameters, denoted as $p(\theta^{*}|\mathcal{D})$, utilizing the posterior distributions of local models, $p(\theta_{k}|\mathcal{D}_{k})$. However, exact posterior inference is typically computationally intractable, often requiring the introduction of approximate inference methods. In this study, we adopt variational learning to approximate local posterior distributions based on a shared prior distribution $p(\theta)$ and client-specific likelihoods $p(\mathcal{D}_{k}|\theta_{k})$. Given a parametric distribution family $\mathcal{Q}$, optimization seeks a distribution $q\in\mathcal{Q}$ that minimizes the KL divergence from the true posterior distribution $p(\theta|\mathcal{D})$, i.e.,
$\min_{q(\theta)\in\mathcal{Q}}\operatorname*{D_{KL}}(q(\theta)\|p(\theta|\mathcal{D})).$ (1)
However, direct optimization of (1) remains challenging, leading to the adoption of the Negative Evidence Lower Bound (ELBO) as a surrogate objective
$\min_{q(\theta)\in\mathcal{Q}}-\mathbb{E}_{q(\theta)}[\log p(\mathcal{D}|\theta)]+\operatorname*{D_{KL}}(q(\theta)\|p(\theta)).$ (2)
We approximate the posterior distributions of the local models $p(\theta_{k}|\mathcal{D}_{k}),~\forall k\in\{1,..,N\}$ by optimizing the objective (2) using the IVON optimizer [43] grounded in the Bayesian learning rule [28]. In contrast to classical variational inference, IVON integrates variational learning directly into the optimization process, without altering the model architecture or loss function. IVON maintains a Gaussian posterior over weights through efficient second-order updates based on reparameterized Hessian estimates, enabling uncertainty-aware learning at a cost comparable to deterministic learning with Adam, thereby making Bayesian deep learning more scalable. Subsequently, these local posteriors are aggregated to obtain the global posterior distribution $p(\theta^{*}|\mathcal{D})$.
Based on this formulation, we introduce the following assumptions regarding the common prior $p(\theta)$ and the parametric family $\mathcal{Q}$, which will hold in our experimental setting presented in Section 5.
Assumption 1 (Mean-field Model).
The parametric family $\mathcal{Q}$ consists of $d$-dimensional Gaussian distributions with independent marginals. Specifically, $\theta\sim\mathcal{N}(\mu,\Sigma)$, where $\mu\in\mathbb{R}^{d}$ is the mean vector and $\Sigma=\operatorname*{diag}(\sigma^{2}_{1},\ldots,\sigma^{2}_{d})$ is a diagonal covariance matrix.
3.2 Aggregation Phase
The aggregation phase refers to the process of combining locally trained model updates from multiple clients into a global model. In the following, we outline and discuss various aggregation techniques for the $N$ local distributions, each consistent with Assumption 1 and parametrized with $\{(\mu_{k},\Sigma_{k})\}_{k=1}^{N}$.
  • Empirical Arithmetic Aggregation employed in [51, 39, 5, 20], also known as naive aggregation, is the weighted average of the distribution statistics:
    $\Sigma_{EAA}=\sum_{k=1}^{N}w_{k}\Sigma_{k},\quad\mu_{EAA}=\sum_{k=1}^{N}w_{k}\mu_{k}.$ (3)
  • Barycentric Aggregation, employed for BFL in [25], uses the barycenter of the clients’ posteriors as an aggregated model, minimizing the average discrepancy across clients. Some examples include:
    • The Wasserstein-2 Barycenter, minimizing the average discrepancy in terms of the Wasserstein-2 distance between the clients’ distributions. A closed form solution for the Wasserstein Barycenter under Assumption 1 is obtained in [2]:
    • The Reverse KL Barycenter, minimizing the average reverse KL divergences between the local distributions, coincides with the multiplicative aggregation of posteriors used in [1, 36, 22, 40]. Its statistics are obtained [31] as follows:
      $\Sigma_{RKL}=\left(\sum_{k=1}^{N}w_{k}\Sigma_{k}^{-1}\right)^{-1},$~ (5)
      $\mu_{RKL}=\Sigma_{RKL}\sum_{k=1}^{N}w_{k}\Sigma_{k}^{-1}\mu_{k}.$ (6)
To support the various aggregation strategies that operate on covariance matrices, and to avoid arbitrarily setting the effective sample size parameter in IVON, which ideally corresponds to the total dataset size, often unknown at the server, we consider a subclass of IVON that explicitly stores the covariance matrix and performs sampling directly from it instead of relying on the Hessian matrix. This is made possible by the relation [43]
$\sigma^{2}=\frac{1}{N(h+\delta)},$ (7)
which enables the covariance matrix to be computed from the Hessian, where $\sigma^{2}$ denotes the variance, $N$ is the dataset size, $h$ is the Hessian approximation, and $\delta$ is the weight decay term. In practice, the covariance matrix is first computed locally after estimating the Hessian. These local covariances are then aggregated on the server, and the resulting global covariance matrix is sent back to the clients, where it is used to reconstruct the updated local Hessians. Unlike previous work [40] that employed IVON but was restricted to RKLB aggregation due to its direct operation on the Hessian, our formulation enables a wider support for covariance-based aggregation strategies.

### v2

@@ Abstract
Bayesian Federated Learning (BFL) combines uncertainty modeling with decentralized training, enabling the development of personalized and reliable models in the presence of data heterogeneity and privacy constraints. Existing approaches typically rely on Markov Chain Monte Carlo (MCMC) sampling or variational inference, often incorporating personalization mechanisms to better adapt to the local data distributions. In this work, we propose an information-geometric projection framework for personalization in parametric BFL. By projecting the global model onto a neighborhood of the user’s local model, our method enables a tunable trade-off between global generalization and local specialization. Under mild assumptions, we show that this projection step is equivalent to computing a barycenter in the statistical manifold, allowing us to derive closed-form solutions and achieve cost-free personalization. We apply the proposed approach within a variational learning setup using the Improved Variational Online Newton (IVON) optimizer and extend it to general aggregation schemes in BFL. Empirical evaluations under heterogeneous data distributions confirm that our method effectively balances global and local performance with minimal computational overhead. Code is available at https://github.com/NourJamoussi/Information-Geometry-for-Bayesian-Federated-Learning.
@@ 1 Introduction
Federated learning (FL) is a collaborative machine learning paradigm designed to preserve data privacy. By connecting a central server to multiple participants, commonly referred to as clients or end-users, FL enables distributed model training while keeping data local to each client. In a typical FL setting, clients train local models on their private datasets, and the server aggregates these locally updated models into a global model. After each round of local updates, clients share their updated model parameters with the server, which aggregates them to refine the global model. The updated global model is then sent back to the clients for further local training McMahan et al. (2017).
@@
In real-world federated learning scenarios, statistical heterogeneity among clients is a pervasive challenge. For instance, variations in Magnetic resonance imaging (MRI) scanners across hospitals can induce feature skewness, where the input data distributions differ substantially between clients. Likewise, label distribution shifts may occur when certain clients have disproportionately more samples from specific classes than others, such as in specialized hospitals focusing on particular diseases compared to general hospitals. This heterogeneity violates the standard assumption of independent and identically distributed (i.i.d.) data, leading to degraded performance and increased uncertainty of the global model when evaluated on local test data.
@@
Personalized Federated Learning (PFL) is a paradigm in which clients adapt the global model to their local data, i.e., perform model personalization, with the goal of mitigating the impact of statistical heterogeneity. In Tan et al. (2023), the authors propose a comprehensive taxonomy of PFL approaches, classifying them into two main categories: global model personalization techniques and personalized model learning techniques. The former category includes methods such as client selection Yang et al. (2021), meta-learning Fallah et al. (2020), and transfer learning Chen et al. (2020), while the latter encompasses strategies such as parameter decoupling Arivazhagan et al. (2019), knowledge distillation Jeong and Kountouris (2023), and clustering-based approaches Sattler et al. (2020). Furthermore, several studies Zhang et al. (2022b); Boroujeni et al. (2024); Kotelevskii et al. (2022); Zhu et al. (2023) have applied Bayesian methods within the PFL framework, demonstrating improved model calibration and uncertainty quantification. The transition from deterministic FL to Bayesian FL (BFL) naturally motivates an investigation of the manifolds to which clients’ and global posterior distributions belong. This perspective enables the exploration of their geometric properties and the definition of meaningful operations on these manifolds, as discussed in Jamoussi et al. (2024).
In this work, we study the PFL paradigm through the lens of Bayesian learning, leveraging principles from Information Geometry to develop a novel personalization method for BFL. Unlike existing approaches, the proposed method does not require fine-tuning, additional training, or access to either local or shared global data, making it a fully private and computationally cost-free personalization technique. By specifying a divergence metric, we obtain personalized model posteriors by projecting the global posterior distribution onto a sphere centered at the local posterior distribution, where the radius encodes the desired degree of personalization specified by the end-user. This formulation enables unrestricted personalization flexibility, allowing users to seamlessly control the trade-off between local adaptation and global generalization without incurring additional computational or privacy costs.
We show that, for any divergence function that is convex in its first argument, the projection problem is equivalent to computing the weighted barycenter between the local and global posterior distributions. This equivalence establishes a conceptual bridge between two fundamental concepts in Information Geometry, geometric projection and geometric barycenter, thereby providing an interpretable relationship between the radius of the projection sphere and the weights in the barycentric aggregation. Consequently, the degree of personalization becomes both intuitive and directly controllable by end-users.
The paper is organized as follows. Section 2 reviews the related work. Section 3 introduces the parametric BFL framework based on variational learning for local training and posterior aggregation to estimate the global posterior model. We extend the approach of Pal et al. (2024) by employing the Improved Variational Online Newton (IVON) optimizer Shen et al. (2024) for multiple aggregation methods. In Section 4, we present a novel personalization strategy for global BFL models using information-geometric projection and show its equivalence to barycentric aggregation for divergence functions convex in their first argument. The proposed method is general and applicable beyond parametric BFL, including non-parametric settings. For the parametric case, we exploit closed-form barycentric solutions for specific divergences, yielding significant computational gains. Section 5 reports the experimental results, comparing aggregation techniques under the variational learning framework and benchmarking our method against state-of-the-art approaches. Finally, Section 6 discusses the broader applicability of the proposed method to continual learning.
2 Background and Related Work
Bayesian Federated Learning.
BFL seeks to integrate the advantages of Bayesian Deep Learning, such as improved model calibration and uncertainty quantification, within the FL framework. Following the classification proposed in Cao et al. (2023), BFL approaches can be broadly categorized into two main types: client-side BFL Zhang et al. (2022b); Zhu et al. (2023); Liu et al. (2023); Boroujeni et al. (2024); Hasan et al. (2024); Bhatt et al. (2024) and server-side BFL Corinzia et al. (2019); Chen and Chao (2020); Al-Shedivat et al. (2020); Guo et al. (2023), with certain overlaps between these groups, reflecting the diversity in model formulations and inference techniques. In Jamoussi et al. (2024), the authors extend the discussion of the parametric client-side BFL framework by categorizing methods according to the techniques used to aggregate local posteriors and propose a unified framework for aggregation rooted in information geometry.
However, within the scope of this work, it is useful to introduce an additional categorization of BFL based on the adopted modeling assumptions, distinguishing between parametric and nonparametric Bayesian methods.
  • Parametric BFL: Parametric Bayesian learning provides a principled framework within Bayesian statistics, wherein the data generation process is assumed to follow a parametric model. Specifically, this approach assumes that the data distribution can be fully characterized by a finite set of parameters. The parametric family most commonly considered in BFL is the Gaussian family, in which both client-specific posteriors and the global model posterior are approximated using Gaussian distributions Zhang et al. (2022b); Ozer et al. (2022); Guo et al. (2023); Kim and Hospedales (2023); Pal et al. (2024); Swaroop et al. (2025). This Gaussian assumption facilitates tractable inference and efficient parameter aggregation.
  • Nonparametric BFL: Unlike parametric BFL, nonparametric Bayesian learning does not restrict the model to a specific family of distributions. Instead, it provides the flexibility to learn directly from the data, allowing the structure and complexity to emerge naturally rather than being predefined. This approach leverages distributions over infinite-dimensional function spaces, such as Gaussian Processes or Dirichlet Processes, as employed in the context of BFL in Yurochkin et al. (2019). Additionally, particle-based variational inference, used in Kassab and Simeone (2022) for BFL, aligns with the nonparametric perspective by using a set of particles to approximate posteriors without assuming a predefined parametric family.
Personalized Bayesian Federated Learning.
Personalized BFL extends the standard BFL framework by incorporating client-specific adaptations into the shared global model. FedPop Kotelevskii et al. (2022) achieves personalization by modeling each client’s data generation process as a combination of fixed shared population parameters, which describe the common data model, and client-specific random effects, which capture heterogeneity in client data, providing personalization and uncertainty estimation capabilities. To efficiently infer these parameters, FedPop employs MCMC methods to approximate local posterior distributions and perform stochastic optimization in a federated setting. Although MCMC enables flexible and asymptotically exact Bayesian inference, it can be computationally expensive in large-scale scenarios. To address this, pFedBayes Zhang et al. (2022b) adopts variational inference to approximate client-specific posteriors, optimizing a tractable Evidence Lower Bound (ELBO) to achieve computational efficiency in BFL. Although pFedBayes presents a promising approach to personalization in BFL, certain aspects of its methodology introduce notable limitations. First, the degree of personalization is uniformly applied across all clients and determined empirically, without an adaptive mechanism or an intuitive rationale for tailoring it to individual client characteristics. Second, it defines the local model as personalized based on the use of variational inference, where the global model serves as a prior within the KL divergence term of the ELBO. This formulation may be viewed as controversial within the Bayesian community, since the classical interpretation of a prior belief assumes independence from the observed data. More recently, pFedVEM Zhu et al. (2023) has leveraged variational inference to estimate the client-specific uncertainty and model deviation by modeling client parameters as Gaussian distributions centered around a global latent variable. It employs a confidence-based aggregation strategy that ensures that clients with lower uncertainty and smaller deviations contribute more to the global model.
Unlike the aforementioned methods, our approach introduces personalization as an additional step independent of the training process, requiring only access to the global and local posteriors, with no need for extra data or fine-tuning. Moreover, our method provides access to three different variants of the model, i.e., local, personalized, and global, as suggested in Divi et al. (2021), with the ability to continually adjust the personalized variant without incurring additional training costs.
Information Geometry and Optimal Transport in FL.
Information geometry Amari (2016) provides a general framework for studying the geometric properties of statistical manifolds through divergence functions, allowing distributions to be intuitively treated as elements of a manifold. This framework naturally extends fundamental geometric concepts, such as projections onto constraint sets Csiszár (1975); Csiszár and Matus (2003) and the identification of barycenters, which serve as statistical centers of mass for a given set of probability distributions Nielsen and Nock (2009); D’Ortenzio et al. (2022).
In contrast, optimal transport Villani (2009) focuses on identifying the most efficient transport plan between two probability distributions with respect to a given cost function. Under specific assumptions on the cost function, optimal transport enables the definition of distance functions on the statistical manifold, as exemplified by the family of Wasserstein-p distances Villani (2009). The Sinkhorn algorithm Cuturi (2013) serves as a natural bridge between information geometry and optimal transport, enabling efficient computation of transport plans by incorporating entropy regularization into the optimal transport problem.
Multi-Marginal Optimal Transport (MMOT) Gangbo and Święch (1998); Pass (2015) extends the classical two-marginal optimal transport problem to multiple distributions by seeking an optimal joint coupling that minimizes a given $n$-ary cost function over all marginal distributions. Furthermore, for specific forms of the $n$-ary cost, the MMOT problem can be shown to be equivalent to computing the Wasserstein barycenter (WB) of the set of marginals. This equivalence highlights MMOT as a more general framework within which the WB emerges as a special case, characterizing the optimal interpolation of multiple distributions in Wasserstein space. Recently, applications of concepts from information geometry and optimal transport have been explored in the FL setting. In Farnia et al. (2022), the authors introduce FedOT, a PFL algorithm that integrates optimal transport with model training. FedOT employs deterministic FL, leveraging MMOT to align data distributions. Specifically, it learns transport maps that transform data points from heterogeneous distributions into a shared domain while simultaneously training a predictive model on the mapped data. Similarly, FedDRO Li et al. (2024) aggregates client data distributions via a Wasserstein barycenter and trains the global model against worst-case perturbations within a Wasserstein ball centered at this barycenter. In the context of BFL, Hassan et al. (2023) considers structured latent variable models in which local latent variables are kept private, formulating a decentralized variational inference problem and proposing a communication-efficient aggregation scheme based on Wasserstein barycenters. Similarly, Jamoussi et al. (2024) introduces a unifying framework for aggregation in BFL grounded in barycentric aggregation.
3 Parametric Bayesian Federated Learning
3.1 Learning Phase
The primary objective of BFL is to estimate the posterior distribution of the global model parameters, denoted as $p(\theta^{*}|\mathcal{D})$, using the posterior distributions of local models, $p(\theta_{k}|\mathcal{D}_{k})$. However, exact posterior inference is typically computationally intractable, necessitating the use of approximate inference methods. In this study, we adopt variational learning to approximate local posterior distributions based on a shared prior distribution $p(\theta)$ and client-specific likelihoods $p(\mathcal{D}_{k}|\theta_{k})$. Given a parametric distribution family $\mathcal{Q}$, optimization seeks a distribution $q\in\mathcal{Q}$ that minimizes the KL divergence from the true posterior distribution $p(\theta|\mathcal{D})$, i.e.,
$\min_{q(\theta)\in\mathcal{Q}}D_{\mathrm{KL}}(q(\theta)\|p(\theta|\mathcal{D})).$ (1)
However, direct optimization of equation 1 is generally intractable, motivating the use of the Negative Evidence Lower Bound (ELBO) as a surrogate objective:
$\min_{q(\theta)\in\mathcal{Q}}-\mathbb{E}_{q(\theta)}[\log p(\mathcal{D}|\theta)]+D_{\mathrm{KL}}(q(\theta)\|p(\theta)).$ (2)
We approximate the posterior distributions of the local models $p(\theta_{k}|\mathcal{D}_{k}),~\forall k\in\{1,\dots,N\}$ by optimizing the objective in equation 2 using the IVON optimizer Shen et al. (2024), which is grounded in the Bayesian learning rule Khan and Rue (2021). Unlike classical variational inference, IVON integrates variational learning directly into the optimization process, without requiring modifications to the model architecture or loss function. IVON maintains a Gaussian posterior over the weights through efficient second-order updates based on reparameterized Hessian estimates, enabling uncertainty-aware learning at a computational cost comparable to deterministic training with Adam. This makes Bayesian deep learning more scalable in practice. Subsequently, the local posteriors are aggregated to obtain the global posterior distribution $p(\theta^{*}|\mathcal{D})$.
Based on this formulation, we introduce the following assumptions regarding the common prior $p(\theta)$ and the variational family $\mathcal{Q}$, which will hold throughout the experimental setting described in Section 5.
Assumption 1 (Mean-field Model).
The variational family $\mathcal{Q}$ consists of $d$-dimensional Gaussian distributions with independent marginals. Specifically, $\theta\sim\mathcal{N}(\mu,\Sigma)$, where $\mu\in\mathbb{R}^{d}$ is the mean vector and $\Sigma=\operatorname*{diag}(\sigma^{2}_{1},\ldots,\sigma^{2}_{d})$ is a diagonal covariance matrix.
3.2 Aggregation Phase
The aggregation phase refers to the process of combining locally trained model updates from multiple clients into a single global model. In what follows, we outline and discuss several aggregation techniques for the $N$ local distributions, each consistent with Assumption 1 and parametrized with $\{(\mu_{k},\Sigma_{k})\}_{k=1}^{N}$.
  • Empirical Arithmetic Aggregation, also known as naive aggregation, is employed in Zhang et al. (2022b); Ozer et al. (2022); Bhatt et al. (2024); Fischer et al. (2024). It computes the weighted average of the distribution statistics:
    $\Sigma_{\mathrm{EAA}}=\sum_{k=1}^{N}w_{k}\Sigma_{k},\quad\mu_{\mathrm{EAA}}=\sum_{k=1}^{N}w_{k}\mu_{k}.$ (3)
  • Barycentric Aggregation, adopted for BFL in Jamoussi et al. (2024), uses the barycenter of the clients’ posteriors as the aggregated model, minimizing the average discrepancy among client distributions. Some examples include:
    • Wasserstein-2 Barycenter: minimizes the average discrepancy in terms of the Wasserstein-2 distance between the clients’ distributions. Under Assumption 1, a closed-form solution for the Wasserstein barycenter is derived in Álvarez-Esteban et al. (2016):
    • Reverse KL Barycenter: minimizes the average reverse KL divergences between the local distributions and coincides with the multiplicative aggregation of posteriors proposed in Al-Shedivat et al. (2020); Liu et al. (2023); Guo et al. (2023); Pal et al. (2024). Its closed-form expressions are given in Koliander et al. (2022):
      $\Sigma_{\mathrm{RKL}}=\left(\sum_{k=1}^{N}w_{k}\Sigma_{k}^{-1}\right)^{-1},\quad\mu_{\mathrm{RKL}}=\Sigma_{\mathrm{RKL}}\sum_{k=1}^{N}w_{k}\Sigma_{k}^{-1}\mu_{k}.$ (5)
To support the different aggregation strategies that operate on covariance matrices, and to avoid arbitrarily setting the effective sample size parameter in IVON, which ideally corresponds to the total dataset size but is often unknown at the server, we adopt a subclass of IVON that explicitly stores the covariance matrix and performs sampling directly from it, rather than relying on the Hessian matrix. This reformulation is made possible by the relation derived in Shen et al. (2024)
$\sigma^{2}=\frac{1}{N(h+\delta)},$ (6)
which links the variance $\sigma^{2}$ to the Hessian approximation hh [^1] For a multivariate Gaussian, the Hessian of the negative log-likelihood is proportional to the inverse covariance matrix. When the covariance is diagonal, the Hessian is also diagonal, so the Hessian approximation reduces to per-parameter scalar entries., the dataset size $N$, and the weight decay term $\delta$. In practice, the covariance matrix is first computed locally after estimating the Hessian. These local covariance matrices are then aggregated at the server, and the resulting global covariance is distributed back to the clients, where it is used to reconstruct the updated local Hessians. Unlike prior work Pal et al. (2024), which employed IVON but was limited to RKLB-based aggregation due to its direct dependence on the Hessian, our formulation enables a broader range of covariance-based aggregation strategies.